OpenStudy (anonymous):

is this correct ? Write the equation of the line passing through the points (0, -4) and (-2, 2).

OpenStudy (anonymous):

m= 6 - (-4)/ -2 - 0 = 10/ -2 = -2 y - (-4) = -2 (x-0) y + 4 = -x + 0 y + 4 - 4 = -x + 0 - 4 y = (-x )-5

OpenStudy (anonymous):

where did the 6 come from?

OpenStudy (abb0t):

Whre is your 2x?

OpenStudy (abb0t):

Yes, also where did the 6 come from?

OpenStudy (anonymous):

srry srry thats for another question

OpenStudy (anonymous):

m= 2 -(-4)/ -2 - 0 = 10/ -2 = -2 y - (-4) = -2 (x-0) y + 4 = -2x + 0 y + 4 - 4 = -2x + 0 - 4 y = (-x )-5

OpenStudy (abb0t):

Where did your -2x go in your final answer? How did you get (-x)?

OpenStudy (anonymous):

m= 2 -(-4)/ -2 - 0 = 10/ -2 = -2 y - (-4) = -2 (x-0) y + 4 = -2x + 0 y + 4 - 4 = -2x + 0 - 4 y = (-2x )-5

OpenStudy (abb0t):

And how did you get -5 if you subtracted 4?

OpenStudy (anonymous):

srry my keyboard is a lil messed up

OpenStudy (anonymous):

@Untamed_Being You need to write down the formula first, then plug the numbers in

OpenStudy (anonymous):

m= 2 -(-4)/ -2 - 0 = 10/ -2 = -2 y - (-4) = -2 (x-0) y + 4 = -2x + 0 y + 4 - 4 = -2x + 0 - 4 y = (-2x )-4

OpenStudy (anonymous):

again... where did you get the 10 from and 10/-2 is not -2

OpenStudy (anonymous):

@Chlorophyll i did

OpenStudy (anonymous):

@completeidiot please leave the convo

OpenStudy (abb0t):

Yes, it should be 6. Not 10.

OpenStudy (anonymous):

Slope formula: m = ( y2 - y1) / ( x2 - x1) = ( 2 + 4) / -2

OpenStudy (anonymous):

m= 2 -(-4)/ -2 - 0 = 6/ -2 = -2

OpenStudy (anonymous):

m = 6/ -2 = .... ?!?

OpenStudy (anonymous):

@Untamed_Being It'll be pointless to go through the whole process when you's already wrong at the first line!

OpenStudy (anonymous):

im following the lesson on my notebook @Chlorophyll

OpenStudy (anonymous):

Plug those values into the point-slope formula, y - y1 = m(x - x1). y - 3 = -1(x - 2) y - 3 = -x + 2 y - 3 + 3 = -x + 2 + 3 y = -x + 5

OpenStudy (anonymous):

Your basic arthmetic is totally off !!! 2 + 4 = 6 ( not 10 !) 6/ -2 = ....?

OpenStudy (anonymous):

i've already said my keybaord is messed up

OpenStudy (anonymous):

Then what's the right value of m now???

OpenStudy (anonymous):

ok this is my question Write the equation of the line passing through the points (0, -4) and (-2, 2). m= 2 -(-4)/ -2 - 0 = 6/ -2 = -2

OpenStudy (anonymous):

is it -2 ?

OpenStudy (anonymous):

So you can't do 6/2 = ?

OpenStudy (anonymous):

thats a fraction though

OpenStudy (anonymous):

6 -- 2

OpenStudy (anonymous):

\[6 \div 2\]

OpenStudy (anonymous):

6 /2 is 3

OpenStudy (anonymous):

ok so then 6/-2 =?

OpenStudy (anonymous):

so it would be m= 2 -(-4)/ -2 - 0 = 6/ -2 = -3

OpenStudy (anonymous):

Now, don't blame your keyboard anymore !!!

OpenStudy (anonymous):

@completeidiot Thank you for offering your shoulder :)

OpenStudy (anonymous):

@Untamed_Being Are you positively m = -3 now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then, proceed to plug it in POINT-SLOPE form:

OpenStudy (anonymous):

y - (-4) = -3 (x-0) y + 4 = -3x + 0 y + 4 - 4 = -3x + 0 - 4 y = (-x )-4

OpenStudy (anonymous):

Yes, to make it less confused: y + 4 = -3x y = -3x - 4

OpenStudy (anonymous):

so i did it correctly

OpenStudy (anonymous):

I don't know why you keep the zero, and also why you drop 3 in front of -x

OpenStudy (anonymous):

less confusing for me that way srry

OpenStudy (anonymous):

y = -3x - 4 ( NOT y = -x - 4)

OpenStudy (anonymous):

y - (-4) = -3 (x-0) y + 4 = -3x + 0 y + 4 - 4 = -3x + 0 - 4 y = -3x - 4

OpenStudy (anonymous):

You can only subtract constant to constant, nothing involve with x term or y term!

OpenStudy (anonymous):

got it